OpenStudy (anonymous):

(g + 4) – 3g = 1 + g

OpenStudy (anonymous):

g=1

OpenStudy (jamesj):

So first (g + 4) - 3g = g - 3g + 4 = -2g + 4 Now that means -2g + 4 = 1 + g Add 2g to both sides .... and ... can you see where to go now?

OpenStudy (anonymous):

yeah. thanks ur a big help.